OpenStudy (anonymous):

5x-8y=3 5y-8x=7 Are they perpendicular?

OpenStudy (anonymous):

write slope intercept from if the produc of 2 equation is -1 ,than they are perpendicular.If the same slope than they are parallel

OpenStudy (anonymous):

5x-8y=3 y =(-5x+3)/-8 y=5/8x -3/8 5y-8x=7 y=(8x+7)/5 y=8/5x+7/5 (5/8)*(8/5)=1 no not perpendicular or paralle either